### v1

# Instability of sphalerons in $\phi^{4}$ models with a false vacuum

Abstract.
A one-parameter family of nonlinear (quartic) Klein-Gordon models having a sphaleron solution is studied. The sphaleron arises from a saddle point between true and false vacua in the energy functional. Its instability is shown be governed by a Heun differential equation after a change of variable. This allows an explicit formulation of the eigenfunctions and eigenvalues to be obtained in terms of local Heun functions. Good approximations are found for certain ranges of the parameter.
1. Introduction
Nonlinear Klein-Gordon models are of widespread interest in mathematics and physics. Various kinds of nonlinear phenomena exist in these models, depending on the nature of the field potential [1, 2, 3].
One very interesting phenomenon is a sphaleron — a static, finite energy, unstable solution having a localized profile [4]. These solutions have been observed numerically in non-symmetric quartic potentials [5]. In general, they arise whenever there is a saddle point between true and false vacua in the energy functional. The field profile is symmetric with a single peak. Asymptotically, the field goes to a background value given by the false vacuum, and the peak of the field is the value of the true vacuum.
Sphalerons exhibit a linear instability, which is easily understood to occur by the following argument about the corresponding eigenfunction equation that governs linear perturbations. Consider the derivative of a sphaleron field. Since it satisfies the linearized field equation, it represents an eigenfunction with zero eigenvalue, called a zero mode. This eigenfunction has a node at the location of the peak of the sphaleron. From a general result in eigenfunction theory (Sturm oscillation theorem), however, the ground-state eigenfunction is always nodeless. Therefore, the ground state must have a negative eigenvalue, which implies the sphaleron is linearly unstable when perturbed by the ground-state mode.
The present paper is devoted to studying this instability mathematically. As main results, it shown that the eigenfunctions and eigenvalues can be obtained through transforming the problem into a Heun equation [6], which is a linear second-order differential equation that generalizes the hypergeometric equation. The eigenfunctions are found to be given by certain local Heun functions. In addition, the discrete spectrum is shown to comprise the negative-eigenvalue ground state, and two positive-eigenvalue modes, in addition to a standard zero mode arising from translation invariance.
As further results, an approximation in terms of elementary functions is found for the ground state eigenvalue and eigenfunction.
The rest of the paper is organized as follows.
In section 2, as a preliminary result, it is shown that a general non-symmetric Klein-Gordon potential with a false vacuum can be transformed into one-parameter family of potentials by a shift, scaling, reflection on the Klein-Gordon field. This family coincides with a potential studied in Ref. [5] and reviewed in Ref. [7]. The sphaleron solution is then introduced, along with the the corresponding moving solitary wave solution obtained by applying a Lorentz boost.
In section 3, the linear perturbation problem is studied for the sphaleron. Its solution in terms of local Heun functions is presented, using a change of variable motivated by one of the singularities in the linear differential equation for the perturbation. The discrete spectrum is determined numerically in conjunction with the Sturm oscillation theorem.
In section 4, explicit approximate expressions for the ground state eigenvalue and eigenfunction are developed in terms of elementary functions.
In section 5, some concluding remarks are given.
An appendix summarizes the conserved quantities of nonlinear Klein-Gordon equations.
2. Sphaleron in a false vacuum
For a general nonlinear KG model of a scalar field $\phi(x,t)$, the equation of motion is given in terms of an interaction potential $V(\phi)$ by
$\phi_{tt}-\phi_{xx}+V^{\prime}(\phi)=0$ (1)
which arises from the action principle
$S[\phi]=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\big{(}{-}\tfrac{1}{2}\phi_{t}^{2}+\tfrac{1}{2}\phi_{x}^{2}+V(\phi)\big{)}dx\,dt.$ (2)
Here the field and the space-time coordinates are taken to be dimensionless (i.e. relativistic units are employed). There is freedom of adding an arbitrary constant $V_{0}$ to the interaction potential $V(\phi)$ without changing the equation of motion. The action principle is invariant under time-translations, space-translations, and Lorentz boosts, which yield respective conservation laws for energy, linear momentum, and boost momentum. See the appendix.
Consider a non-symmetric quartic potential with a false vacuum. This means that $V(\phi)$ has a local maximum and two local minimums, one of which is a global minimum. In physical terms, the global minimum represents the true vacuum, while the other minimum represents the false vacuum. These conditions require, firstly, that $V^{\prime}(\phi)$ must have three distinct real roots: $\phi=a_{1},a_{2},a_{3}$ where, without loss of generality, $a_{1}<a_{2}<a_{3}$. Secondly, $V^{\prime}(\phi)$ must not be anti-symmetric when reflected around $\phi=a_{2}$: $a_{1}-a_{2}\neq a_{3}-a_{2}$. Thus, we must have $V^{\prime}(\phi)=\alpha(\phi-a_{1})(\phi-a_{2})(\phi-a_{3})$, with $a_{1}<a_{2}<a_{3}$ and $a_{2}\neq\tfrac{1}{2}(a_{3}+a_{1})$, where $\alpha>0$.
[…]
$V(\phi)=\alpha\big{(}\tfrac{1}{4}\phi^{4}-\tfrac{1}{3}\tilde{a}\phi^{3}-\tfrac{1}{2}\tilde{b}^{2}\phi^{2}+\tilde{a}\tilde{b}^{2}\phi\big{)}+V_{0}$ (3)
whose local extrema are given by $V(\tilde{a})=V_{0}+\tfrac{1}{12}\alpha(6\tilde{a}_{1}^{2}-\tilde{a}_{2}^{2})\tilde{a}_{2}^{2}$ where $\phi=\tilde{a}$ is the maximum, and $V(\pm\tilde{b})=V_{0}+\tfrac{1}{12}\alpha(\tilde{a}\pm 3\tilde{b})\tilde{b}^{3}$ where $\phi=\pm\tilde{b}$ are the two minimums. The difference in the potential at these minima, which distinguishes the true vacuum from the false vacuum, is given by
$\Delta V:=V(\tilde{b})-V(-\tilde{b})=\tfrac{4}{3}\alpha\tilde{a}\tilde{b}^{3}.$ (4)
There are two cases to consider, depending on sign of $\tilde{a}\neq 0$.
[…]
Case $\tilde{a}>0$: the vacua are interchanged, with $\phi=\tilde{b}$ being the false vacuum and $\phi=-\tilde{b}$ the true vacuum, where $\tilde{b}<0<\tilde{a}<\tilde{b}$. The potential (3) in this case supports a solitary wave on a bright background, with $\phi\to\tilde{b}$ at $x\to\pm\infty$. This wave describes a ground state lump when it is stationary. Using similar steps as in the previous case, we can shift the potential to get
$)V(\phi)=\alpha\phi^{2}\big{(}\tfrac{1}{4}\phi^{2}+(\tfrac{1}{3}\tilde{a}-\tilde{b})\phi+\tilde{b}(\tilde{b}-\tilde{a}$$\big{)}$ (6)
[…]
It will now be useful to express the potentials (5) and (6) in a simpler factorized form by a reparameterization. This also leads to a very simple expression for the ground state lumps. Hereafter we will focus on the potential (5) in which the ground state lump is bright. The opposite case in which the ground state lump is dark is given by a reflection $\phi\to-\phi$, corresponding to the potential (6).
The reparameterization of the potential (5) is achieved by expressing the pair of non-zero roots as
$\tfrac{2}{3}\big{(}3\tilde{b}-|\tilde{a}|-\sqrt{|\tilde{a}|(3\tilde{b}+|\tilde{a}|)}\big{)}=b\tanh(a),\quad\tfrac{2}{3}\big{(}3\tilde{b}-|\tilde{a}|+\sqrt{|\tilde{a}|(3\tilde{b}+|\tilde{a}|)}\big{)}=b{\rm coth}(a)$ (7)
where $a>0$ and $b>0$. Equivalently, we have
$).b=2\sqrt{\tilde{b}(\tilde{b}-\tilde{a})},\quad a=\tfrac{1}{2}{\rm arctanh}\Big{(}\sqrt{1-\tilde{a}/\tilde{b}}\big{/}(1-\tfrac{1}{3}\tilde{a}/\tilde{b}$$\Big{)}$. (8)
A scaling transformation $\phi\to b\phi$ combined with a dilation $(x,t)\to\sqrt{\tfrac{2}{\alpha}}(x,t)$ then leads to the following characterization of the resulting potential.
Proposition 1.
Up to a shift, scaling, and reflection on $\phi$, and a dilation on $(t,x)$, any non-symmetric quartic potential with a false vacuum belongs to the $1$-parameter family
$V(\phi)=2\phi^{2}(\phi-\tanh(a))(\phi-{\rm coth}(a)),\quad a>0.$ (9)
The false vacuum is $V=0$ at $\phi=0$, and the true vacuum has $V<0$ at
$\phi=\tfrac{3}{4}{\rm coth}(2a)\Big{(}1+\sqrt{1-\tfrac{8}{9}\tanh(2a)^{2}}\Big{)}$ (10)
which lies between $\phi=\tanh(a)$ and $\phi={\rm coth}(a)$. The equation of motion (1) in this potential is given by
$\phi_{tt}-\phi_{xx}+4(2\phi^{2}-3{\rm coth}(2a)\phi+1)\phi=0.$ (11)
This family (9) is equivalent, up to scaling, to the potential
[…]
with $1>\tilde{b}>0$, $\tilde{a}>0$ studied in Ref. [5], and also to the potential
$V(\phi)=4\phi^{2}(1-\phi/\tilde{b}^{2})(1-\phi/\tilde{a}^{2})$ (13)
with $\tilde{b}>\tilde{a}>0$ listed in Ref. [7]. A plot is shown in Fig. 1.
Figure 1. Potential with a false vacuum at $\phi=0$: $a=$ 0.5, 0.8, 1.5
2.1. Ground states
A ground state of a general nonlinear KG model (1) is a static solution $\phi=\phi(x)$ satisfying the ODE
$\phi_{xx}=V^{\prime}(\phi).$ (14)
This is an oscillator equation with effective potential $V_{\text{eff.}}(\phi)=-V(\phi)$. Integration gives
$\tfrac{1}{2}\phi_{x}^{2}-V(\phi)=E_{0}$ (15)
where $E_{0}=\text{const.}$ is the oscillator energy.
The shape of the effective potential determines the allowed non-constant ground states $\phi(x)$. For a non-symmetric potential (9), the effective oscillator potential is an inverted quartic
$V_{\text{eff.}}(\phi)=-2\phi^{2}(\phi-\tanh(a))(\phi-{\rm coth}(a))$ (16)
which has a local maximum at $\phi=0$ and a global maximum at
$\phi=\tfrac{3}{4}{\rm coth}(2a)\big{(}1+\sqrt{1-\tfrac{8}{9}\tanh(2a)^{2}}\big{)},$ (17)
in addition to a local minimum at
$\phi=\tfrac{3}{4}{\rm coth}(2a)\big{(}1-\sqrt{1-\tfrac{8}{9}\tanh(2a)^{2}}\big{)}>0.$ (18)
It is straightforward to see that the only non-constant ground states supported in this potential (16) consist of a bright lump where $\phi$ goes from 0 at $x=\pm\infty$ to a peak $\phi_{\max}=\tanh(a)$ at some position $x=x_{0}$. This lump thereby has $E_{0}=0$. Integration of the oscillator equation (15) yields the explicit expression
$\phi(x)=\sinh(2a)/\big{(}\cosh(2a)+\cosh(2(x-x_{0}))\big{)}.$ (19)
This solution was first derived in a slightly different form for the equivalent potentials (12) and (13). In particular, $\phi(x)=\tilde{b}/\big{(}1+(1-\tilde{a})\sinh(\tilde{b}x/\sqrt{\tilde{a}})^{2}\big{)}$ and $\phi(x)=\tilde{a}^{2}\tilde{b}^{2}/\big{(}\tilde{a}^{2}+(\tilde{b}^{2}-\tilde{a}^{2})\cosh(x)^{2}\big{)}$ appear respectively in Ref. [5] and Ref. [7]. The latter also states another very useful form of the solution, which is given by
$\phi(x)=\tfrac{1}{2}\big{(}\tanh(x-x_{0}+a)-\tanh(x-x_{0}-a)\big{)}$ (20)
for the potential (9). This shows that the lump (19) can be viewed as the superposition of a kink and an antikink. It has a peak height
$\phi(x_{0})=\tanh(a)$ (21)
and its full-width is
$\Delta x={\rm arccosh}\big{(}2\cosh(2a)+\sqrt{3\cosh(2a)^{2}+6}\big{)}$ (22)
as defined by where the convexity of $\phi(x)$ is maximum. See Fig. 2.
Figure 2. Lump solution (19) in the false-vacuum potential for $a=$ 5.5, 2.5, 1.0, 0.5
The energy of a ground state is given by the integral
$E[\phi]=\int_{-\infty}^{\infty}\Big{(}\tfrac{1}{2}\phi_{x}^{2}+V(\phi)\Big{)}\,dx=2\int_{-\infty}^{\infty}V(\phi)\,dx.$ (23)
Since the lump (19) has a single peak and decays to zero, its energy can be expressed as a Bogomolny integral
$E[\phi]=2\int_{0}^{\phi_{\max}}\sqrt{2V(\phi)}\,d\phi$ (24)
where $\phi_{\max}=\tanh(a)$. This yields
$E_{\text{lump}}=\tfrac{2}{3}+2\big{(}1-2a{\rm coth}(2a)\big{)}/\sinh(2a)^{2}$ (25)
for the lump energy. (A more complicated expression for the energy is given in Ref. [7] using the alternative form (13) for the potential.) Because the lump is stationary, it clearly has zero momentum, $P_{\text{lump}}=0$, and zero boost-momentum, $J_{\text{lump}}=0$.
2.2. Travelling waves
A Lorentz boost can be applied to any static solution $\phi=\phi(x)$ of the equation of motion, yielding a travelling wave solution $\phi=\phi((x-\nu t)/\sqrt{1-\nu^{2}})$ with speed $\nu$. Travelling waves move rightward if $\nu>0$, or leftward if $\nu<0$. A physical requirement is that $|\nu|<1$, so their motion lies inside the light cone; solutions with $|\nu|>1$ are called tachyonic.
The ground state lump (19) gives rise to solitary waves
$\phi(x,t)=\sinh(2a)/\big{(}\cosh(2a)+\cosh(2(\xi-\xi_{0}))\big{)}$ (26)
in terms of travelling wave variable
$\xi=\frac{x-\nu t}{\sqrt{1-\nu^{2}}},\quad|\nu|<1.$ (27)
An equivalent expression for the solitary waves is given by
$\phi(x,t)=\tfrac{1}{2}\big{(}\tanh(\xi-\xi_{0}+a)-\tanh(\xi-\xi_{0}-a)\big{)}.$ (28)
This describes the superposition of a kink and an antikink which move in the same direction.
The solitary wave has energy
$E_{\text{solitary}}=\frac{E_{\text{lump}}}{\sqrt{1-\nu^{2}}}=\frac{2}{\sqrt{1-\nu^{2}}}\Big{(}\tfrac{1}{3}+\big{(}1-2a{\rm coth}(2a)\big{)}/\sinh(2a)^{2}\Big{)}$ (29)
and momentum
$P_{\text{solitary}}=\frac{\nu E_{\text{lump}}}{\sqrt{1-\nu^{2}}}=\frac{2\nu}{\sqrt{1-\nu^{2}}}\Big{(}\tfrac{1}{3}+\big{(}1-2a{\rm coth}(2a)\big{)}/\sinh(2a)^{2}\Big{)}.$ (30)
These expressions give an energy-momentum 2-vector $(E_{\text{solitary}},P_{\text{solitary}})$ which is the boost of the lump 2-vector $(E_{\text{lump}},0)$.
3. Linear Instability
To investigate linear stability of a ground state $\phi=\phi(x)$, consider a small perturbation with frequency $\omega$,
$\varphi(x,t)=\phi(x)+\epsilon\eta(x)\exp(i\omega t)$ (31)
where $|\epsilon|\ll 1$. Linearization of the equation of motion for $\varphi(x,t)$ around $\epsilon=0$ yields
$0=\varphi_{tt}-\varphi_{xx}+V^{\prime}(\varphi)=\epsilon\exp(i\omega t)\Big{(}{-}\omega^{2}\eta-\eta_{xx}+V^{\prime\prime}(\phi(x))\eta)\Big{)}+O(\epsilon^{2}),$ (32)
which gives a linear Schrödinger equation for $\eta(x)$:
$\eta_{xx}+U(x)\eta=\omega^{2}\eta$ (33)
with the potential
$U(x)=V^{\prime\prime}(\phi(x)).$ (34)
This is an eigenfunction problem for the linear operator $\partial_{x}^{2}+U(x)$, with eigenvalue $\lambda=\omega^{2}$.
If all eigenvalues are non-negative, then $\omega$ is real and thus the perturbation (31) is oscillatory in $t$, whereby the ground state $\phi(x)$ would be linearly stable. A zero eigenvalue, $\lambda_{0}=0$, is given by the eigenfunction $\eta(x)=\phi^{\prime}(x)$, which is called the translational mode or zero mode. If at least one negative eigenvalue exists, $\lambda_{-1}<0$, it gives $\omega=\pm i\sqrt{|\lambda_{-1}|}$, which implies the perturbation (31) is exponential in time, since $\exp(i\omega t)=\exp(\pm\sqrt{|\lambda_{-1}|}\,t)$. Existence of the growing mode $\eta(x)\exp(\sqrt{|\lambda_{-1}|})$ would indicate that the ground state $\phi(x)$ is linearly unstable.
The goal is now to prove rigorously that the ground state lump (19) in the false vacuum of the potential (9) is linearly unstable.
3.1. Spectrum of the perturbation potential
The potential (9) has $V^{\prime\prime}(\phi(x))=4(6\phi(x)^{2}-3(\tanh(a)+{\rm coth}(a))\phi(x)+1)$. For the ground state lump (19), this yields the perturbation potential
$U(x)=4-24\frac{1+\cosh(2x)\cosh(2a)}{(\cosh(2x)+\cosh(2a))^{2}}$ (35)
which has the following main features. It is asymptotically constant, with $U(x)\to 4$ exponentially as $|x|\to\infty$, while $U(x)<4$ for finite $x$. It is symmetric in $x$, such that $x=0$ is the minimum when ${\rm coth}(a)\geq\sqrt{3}$ or the local maximum when ${\rm coth}(a)<\sqrt{3}$, thus giving a double-well in the latter case and a single-well in the former case. See Fig. 3. In addition, the zero mode of this potential is given by
$\eta(x)=\phi^{\prime}(x)=-2\sinh(2a)\sinh(2x)/(\cosh(2x)+\cosh(2a))^{2},\quad\lambda_{0}=0$ (36)
which has a single node, $\eta=0$ at $x=0$.
Figure 3. Perturbation potential: $a=$ 0.25, ${\rm arctanh}\tfrac{1}{\sqrt{3}}$ (=0.658), 1.5
These features allow the spectrum of $\partial_{x}^{2}+U(x)$ to be deduced from basic results about linear Schrödinger operators (see e.g. Ref. [8]). Firstly, the continuous spectrum consists of $4,\infty)$, by Weyl’s theorem. Secondly, the lowest eigenvalue in the discrete spectrum will be negative, $\lambda_{-1}<0$, corresponding to a nodeless eigenfunction $\eta_{-1}(x)$, due to Sturm’s oscillation theorem.
Therefore, the ground state lump (19) is linearly unstable and thereby it is a sphaleron. Remarkably, the eigenfunction problem
$\eta_{xx}+U(x)\eta=\lambda\eta,\quad\lambda=\omega^{2}$ (37)
can be solved in terms of Heun functions, as will be shown next through an explicit transformation. This will lead to an analytical way to find the negative eigenvalue $\lambda_{-1}$.
To begin, we will examine the singular points of the linear differential equation (37). The potential (35) is singular at the finite points $x_{0}=\pm a+i\pi/2$ (modulo $\pi$), and each of them is a regular singular point, since $(x-x_{0})^{2}U(x)$ is analytic around $x=x_{0}$. Furthermore, the point at infinity, $x=\infty$, is irregular singular, since the potential is not analytic in $1/x$. We can transform this point into a regular singular point by a change of variable of the form $x\to y=f(\cosh(x))$ where $f$ is some rational function.
The most effective choice turns out to be given by the change of variable
$y=1/\cosh^{2}(x)$ (38)
under which $U(x)$ becomes the following rational function of $y$:
$U=4+\frac{12y(\sinh^{2}(a)y-\cosh(2a))}{(\sinh^{2}(a)y+1)^{2}}.$ (39)
Since $\dfrac{dy}{dx}=-2y\sqrt{1-y}$, the eigenfunction equation (37) is correspondingly transformed to the rational form
$\tilde{\eta}^{\prime\prime}(y)+\frac{3y-2}{2y(y-1)}\tilde{\eta}^{\prime}(y)+\Big{(}\frac{4-\lambda}{4y^{2}(y-1)}+\frac{3(\sinh^{2}(a)y-\cosh(2a))}{y(y-1)(\sinh^{2}(a)y+1)^{2}}\Big{)}\tilde{\eta}(y)=0$ (40)
for the function
$\tilde{\eta}(y)=\eta(x)=\eta({\rm arccosh}(1/\sqrt{y})).$ (41)
We see that coefficients of $\tilde{\eta}^{\prime}(y)$ and $\tilde{\eta}(y)$ are singular at $y=0,1,-1/\sinh^{2}(a)$ and also at $y=\infty$. These are readily checked to be regular singular points of the linear ODE (40).
Now recall that any second-order linear ODE with four regular singular points is equivalent under a change of variables to Heun’s ODE [6]
$H^{\prime\prime}(z)+\big{(}\gamma/z+\delta/(z-1)+\epsilon/(z-p)\big{)}H^{\prime}(z)+\big{(}(\alpha\beta z-q)/(z(z-1)(z-p)\big{)}H(z)=0$ (42)
with $\epsilon=\alpha+\beta-\gamma-\delta+1$, which has regular singular points $z=0,1,p,\infty$, where $\alpha$, $\beta$, $\gamma$, $\delta$, and $q$ are constant parameters. Solutions $H(z)$ are given by an analytic Frobenius series around any one of the points $z=0,1,p$, in terms of the local Heun function $H\ell(p,q,\alpha,\beta,\gamma,\delta;z)$ which is analytic in $z$ with $H\ell(p,q,\alpha,\beta,\gamma,\delta;0)=1$. The main properties of local Heun functions are summarized in Ref. [6].
The explicit transformation of the linear ODE (40) to Heun’s ODE (42) is given by
$H(y)=\frac{\tilde{\eta}(y)}{y^{\sqrt{1-\lambda/4}}(\sinh(a)^{2}y+1)^{3}}$ (43)
with
$\alpha=\tfrac{1}{2}\sqrt{4-\lambda}+3,\quad\beta=\tfrac{1}{2}(\sqrt{4-\lambda}+7),\quad\gamma=\sqrt{4-\lambda}+1,\quad\delta=\tfrac{1}{2},\quad\epsilon=6$ (44)
and
$p=-\frac{1}{\sinh^{2}(a)},\quad q=\frac{6\cosh(2a)(\sqrt{4-\lambda}+3)-(\sqrt{4-\lambda}+6)(\sqrt{4-\lambda}+1)}{4\sinh^{2}(a)}.$ (45)
In terms of $x$, we have
$\eta(x)={\rm sech}(x)^{\sqrt{4-\lambda}}(\sinh(a)^{2}{\rm sech}^{2}(x)+1)^{3}H({\rm sech}^{2}(x)),$ (46)
where the finite singular points $y=0,1,-1/\sinh^{2}(a)$ respectively correspond to $x=\pm\infty,0,\pm a+i\pi/2$ (modulo $\pi$).
From the general features of the potential (35), we can deduce that the eigenfunction $\eta_{-1}(x)$ with the negative eigenvalue $\lambda_{-1}$ must be analytic in $x$, and decay like $e^{-\sqrt{4+|\lambda_{-1}|}\,|x|}$ as $|x|\to\infty$, in addition to being symmetric around $x=0$ and nodeless. In particular, as a consequence of analyticity, the symmetry $\eta_{-1}(-x)=\eta_{-1}(x)$ will hold if and only if $\eta_{-1}^{\prime}(0)=0$. Through expression (43) combined with the transformation (38), the preceding properties imply that
(i) $H(y)$ must be analytic in $y$, non-singular at $y=0$, non-zero for $0<y<1$;
(ii) $\sqrt{1-y}\big{(}y^{\sqrt{1-\lambda/4}}(\sinh(a)^{2}y+1)^{3}H(y)\big{)}^{\prime}$ must vanish at $y=1$.
Condition (i) implies that, after normalization $H(0)=1$, the function $H(y)$ must be equal to the local Heun function:
$H(y)=H\ell(p,q;\alpha,\beta,\gamma,\delta;y).$ (47)
Around $y=1$, this function (47) has a Frobenius series expansion in 1y1-y with indicial exponents 0 and $1-\delta=1/2$ (using parameters (44)):
$H(y)=$ $C_{1}H\ell(1-p,\alpha\beta-q;\alpha,\beta,\delta,\gamma;1-y)$ (48)
$+C_{2}\sqrt{1-y}H\ell(1-p,\alpha^{\prime}\beta^{\prime}-q^{\prime};\alpha^{\prime},\beta^{\prime},\delta^{\prime},\gamma;1-y)$
where
$q^{\prime}=q+p\gamma(1-\delta),\quad\alpha^{\prime}=\alpha-\delta+1,\quad\beta^{\prime}=\beta-\delta+1,\quad\delta^{\prime}=2-\delta.$ (49)
Thus, $H(1)=H\ell(p,q;\alpha,\beta,\gamma,\delta;1)$ is a finite value, due to $H\ell(1-p,\alpha\beta-q;\alpha,\beta,\delta,\gamma;0)=H\ell(1-p,\alpha^{\prime}\beta^{\prime}-q^{\prime};\alpha^{\prime},\beta^{\prime},\delta^{\prime},\gamma;0)=1$. Condition (ii) thereby yields, after simplification,
$\lim_{y\to 1}\sqrt{1-y}H\ell(p,q;\alpha,\beta,\gamma,\delta;y)^{\prime}=0$ (50)
which constitutes the determining equation for the negative eigenvalue $\lambda=\lambda_{-1}$. This equation can be expressed in a useful equivalent form as follows. First, via substituting the relation (48) and using the finiteness of $H(1)$, we obtain $C_{2}=0$. Next, we evaluate relation (48) for $y=0$ and $y=1$, which gives
$C_{1}$ $=H\ell(p,q;\alpha,\beta,\gamma,\delta;1)$ (51)
$C_{2}$ $=\big{(}1-C_{1}H\ell(1-p,\alpha\beta-q;\alpha,\beta,\delta,\gamma;1)\big{)}/H\ell(1-p,\alpha^{\prime}\beta^{\prime}-q^{\prime};\alpha,^{\prime}\beta^{\prime},\delta^{\prime},\gamma;1)$
where
$\begin{gathered}q^{\prime}=\frac{6\cosh(2a)(\sqrt{4-\lambda}+3)-(\sqrt{4-\lambda}+8)(\sqrt{4-\lambda}+1)}{4\sinh^{2}(a)},\\ \alpha^{\prime}=\tfrac{1}{2}(\sqrt{4-\lambda}+7),\quad\beta^{\prime}=\tfrac{1}{2}\sqrt{4-\lambda}+4,\quad\delta^{\prime}=\tfrac{3}{2}.\end{gathered}$ (52)
Finally, the denominator in expression (51) for $C_{2}$ turns out to be finite and non-zero. Hence, the determining equation $C_{2}=0$ becomes
$H\ell(p,q;\alpha,\beta,\gamma,\delta;1)H\ell(1-p,\alpha\beta-q;\alpha,\beta,\delta,\gamma;1)-1=0$ (53)
in terms of the parameters (44). This establishes the first main result.
Theorem 1.
The ground state eigenvalue $\lambda_{-1}$ of the potential (35) satisfies equation (53) in terms of the local Heun function $H\ell(p,q,\alpha,\beta,\gamma,\delta;z)$. For any fixed value of the parameter $a$ in the potential, this equation (53) can be solved numerically. A plot of $\lambda_{-1}$ as a function of $a$ is shown in Fig. 4.
From a numerical plot, we see that there is a single negative solution $\lambda_{-1}$ for each value of the parameter $a$. The resulting solution of the eigenfunction equation (40) is given by
$\tilde{\eta}_{-1}(y)=y^{\sqrt{1-\lambda_{-1}/4}}(\sinh(a)^{2}y+1)^{3}H\ell(p,q;\alpha,\beta,\gamma,\delta;y)\big{|}_{\lambda=\lambda_{-1}}.$ (54)
Figure 4. Ground state eigenvalue: $\lambda_{-1}$
For numerical purposes, we note that the power series in $y$ which defines the Heun function (47) has radius of convergence $\min(1,|p|)$, where $|p|=1/\sinh(a)^{2}<1$ when $a>{\rm arctanh}(\tfrac{1}{\sqrt{2}})$. Consequently, for an arbitrary value of $a$, the series does not cover the whole interval $0<y\leq 1$. However, from the relation (48) with $C_{2}=0$, we see that the Heun function
$H(y)=H\ell(1-p,\alpha\beta-q;\alpha,\beta,\delta,\gamma;1-y)$ (55)
differs only by the factor $C_{1}$ while the radius of convergence of its defining power series in 1y1-y is $\min(1,|1-p|)=1$ since $1-p=1/\tanh(a)^{2}\geq 1$. Thus, for any value of $a$, the whole interval $0<y\leq 1$ is covered. Using this function, we obtain the equivalent (up to normalization) eigenfunction
$\tilde{\eta}_{-1}(y)=y^{\sqrt{1-\lambda_{-1}/4}}(\sinh(a)^{2}y+1)^{3}H\ell(1-p,\alpha\beta-q;\alpha,\beta,\delta,\gamma;1-y)\big{|}_{\lambda=\lambda_{-1}}.$ (56)
A plot is shown Fig. 5.
Figure 5. $\tilde{\eta}_{-1}(y)$ for $a=$ 2.65 (space dot), 2.14 (dot), 1.44 (dash dot), 1.00 (solid), 0.881 (dash), 0.647 (space dash), 0.100 (long dash)
The corresponding eigenfunction is given by expressions (41), (43), (55), which leads to the following main result.
Theorem 2.
The ground state eigenfunction has the form
$\eta_{-1}(x)={\rm sech}(x)^{\sqrt{4-\lambda_{-1}}}(\sinh(a)^{2}{\rm sech}^{2}(x)+1)^{3}H\ell(1-p,\alpha\beta-q;\alpha,\beta,\delta,\gamma;\tanh^{2}(x))\big{|}_{\lambda=\lambda_{-1}}$ (57)
in terms of the parameters (44) and the local Heun function $H\ell$. A numerical plot of the normalized ground state is shown in Fig. 6 for various values of the parameter $a$ in the potential (35).
These eigenfunctions have a single peak, at $x=0$, for $a\gtrsim 1.00$, and a double peak for $a\lesssim 1.00$.
Figure 6. Ground-state eigenfunction, normalized in $L^{2}$, for $a=$ 2.65 (space dot), 2.14 (dot), 1.44 (dash dot), 1.00 (solid), 0.881 (dash), 0.647 (space dash), 0.100 (long dash)
The eigenvalue equation (53) can be investigated numerically to look for other solutions, which would describe symmetric eigenfunctions. We find that it possesses one positive solution, which we denote as $\lambda_{1}$. Sturm’s oscillation theorem implies that the corresponding symmetric eigenfunction, $\eta_{1}(x)=\tilde{\eta}_{1}(y)$, has two nodes. It is called an internal vibrational (or shape) mode. Fig. 7 shows a plot of the eigenvalue and corresponding eigenfunction.
Figure 7. (Left) $\lambda_{1}$; (Right) $\eta_{1}(x)$, normalized in $L^{2}$, for $a$ = 0.695, 1.082, 1.771, 2.052, 2.555, 4.218
Since the continuous spectrum starts at $\lambda=4$, what remains to be determined is if there is another positive eigenvalue $\lambda_{2}>\lambda_{1}$ corresponding to an eigenfunction that has three nodes. The determining condition for the eigenvalue is simply that $y^{\sqrt{1-\lambda/4}}(\sinh(a)^{2}y+1)^{3}H(y)$ must vanish at $y=1$. We find that it does have a solution, which is plotted in Fig. 8. The corresponding eigenfunction $\eta_{2}(x)=\tilde{\eta}_{2}(y)$ is also plotted. This represents an additional internal vibrational mode.
Figure 8. (Left) $\lambda_{2}$; (Right) $\eta_{2}(x)$, normalized in $L^{2}$, for $a$ = 0.695, 1.656, 2.649, 3.438, 4.218
This completes the determination of the spectrum.
4. Approximate analytical formula for the unstable mode
No closed form expression is known for the local Heun functions appearing in the ground state eigenvalue equation (53) and in the ground state eigenfunction (57). However, the eigenvalue $\lambda_{-1}$ can be shown to satisfy an explicit continued-fraction equation arising from expressions (47)–(48) as follows.
Since the eigenvalue equation is equivalent to the condition $C_{2}=0$, expressions (47)–(48) with $\lambda=\lambda_{-1}$ thereby show that the functions $H\ell(p,q;\alpha,\beta,\gamma,\delta;y)|_{\lambda=\lambda_{-1}}$ and $H\ell(1-p,\alpha\beta-q;\alpha,\beta,\delta,\gamma;1-y)|_{\lambda=\lambda_{-1}}$ are analytic at both $y=0$ and $y=1$. As known from general results stated in Ref. [6], this analyticity is equivalent to a condition on the coefficients in the corresponding series expansions. We will use the series for $H\ell(1-p,\alpha\beta-q;\alpha,\beta,\delta,\gamma;1-y)|_{\lambda=\lambda_{-1}}$ because, for any value of $a$, its convergence interval includes both points $y=0$ and $y=1$. The series coefficients, $c_{i}$, satisfy a three term recurrence relation $R_{j}c_{j+1}-(Q_{j}+\alpha\beta-q)c_{j}+P_{j}c_{j-1}=0$, $j=1,2,\ldots$, with $c_{0}=0$ and $c_{1}=1$, where
$R_{j}=(1-p)(j+1)(j+\delta),\quad Q_{j}=j((j-1+\delta)(2-p)+(1-p)\gamma+\epsilon),\quad P_{j}=(j-1+\alpha)(j-1+\beta).$ (58)
The condition of analyticity at the two points $y=0$ and $y=1$ is given by the continued fraction equation [6]
$q=\frac{R_{0}P_{1}}{Q_{1}+q-}\cdots\frac{R_{j}P_{j+1}}{Q_{j+1}+q-}\cdots.$ (59)
This becomes an explicit equation once the parameters (44) have been substituted. It can be used to get an approximation for $\lambda=\lambda_{-1}$ by truncating the expansion. Truncation at $j=0$ gives a quartic equation in terms of $\mu=\sqrt{4-\lambda_{-1}}$,
[…]
which yields an explicit approximate formula for the eigenvalue $\lambda_{-1}$. Compared to the actual numerical value, this approximation turns out to be good only in a small parameter range $a\lesssim 1$.
An approximate asymptotic formula for both the eigenvalue and eigenfunction can be obtained near the end points $a\to 0$ and $a\to\infty$ of the parameter range. From Fig. 4, note that $\lambda_{-1}\to-5$ as $a\to 0$, and $\lambda_{-1}\to 0$ as $a\to\infty$.
For $a$ near 0, the approximate equation (60) is applicable and yields $\mu\simeq 3-\tfrac{8}{7}a^{2}$ to lowest order. This gives
$\lambda_{-1}\simeq-5+\tfrac{48}{7}a^{2}.$ (61)
Substituting this expression into the eigenfunction equation (40) leads to the solution
$\tilde{\eta}_{-1}(y)=y^{\frac{3}{2}-\frac{4}{7}a^{2}}H\ell_{\text{c}}(0,3-\tfrac{8}{7}a^{2},-1/2,9a^{2},-\tfrac{54}{7}a^{2}-1/2,y)$ (62)
where $H\ell_{\text{c}}$ denotes the confluent local Heun function [6]. Plotting this function shows that it can be well approximated by $H\ell_{\text{c}}(0,3-\tfrac{8}{7}a^{2},-1/2,9a^{2},-\tfrac{54}{7}a^{2}-1/2,y)\simeq 1-2a^{2}y$. The transformation (38) then yields $\eta_{-1}(x)\simeq(1+2a^{2}{\rm sech}(x)^{2}){\rm sech}(x)^{3+\frac{8}{7}a^{2}}$. After expanding ${\rm sech}(x)^{a^{2}}\simeq 1+a^{2}\ln{\rm sech}(x)$, followed by normalizing so that $\eta_{-1}(0)=1$, we finally obtain the approximate ground state eigenfunction
$\eta_{-1}(x)\simeq{\rm sech}(x)^{3}\big{(}1+\tfrac{2}{7}a^{2}\big{(}7\tanh(x)^{2}-4\ln({\rm sech}(x))\big{)}\big{)}$ (63)
when $a$ is near 0.
For large $a$, there is no a priori approximation available for the eigenvalue. Instead we start from the exact eigenfunction (54), put $1-\lambda_{-1}/4=(1+\mu)^{2}$ with $\mu$ assumed to be small, and approximate the parameters (44):
$p\simeq-4e^{-2a},\quad q\simeq 15+6\mu+6e^{-2a},\quad\alpha\simeq 4+\mu,\quad\beta\simeq\tfrac{9}{2}+\mu,\quad\gamma\simeq 3+2\mu,\quad\delta=1/2.$ (64)
We next examine the eigenvalue equation (53) numerically and find that it yields $\mu\simeq 12e^{-4a}$. The eigenvalue is thus given by
$\lambda_{-1}\simeq-96e^{-4a}.$ (65)
This gives the following approximation for the eigenfunction
$\tilde{\eta}_{-1}(y)\simeq\tfrac{1}{8}y^{1+12e^{-4a}}(4+ye^{2a})^{3}H(y)$ (66)
where
$H(y)=H\ell(-4e^{-2a},6e^{-2a}+15,12e^{-4a}+4,\tfrac{9}{2}+12e^{-4a},24e^{-4a}+3,1/2,y).$ (67)
Now we would like to approximate this function in terms of elementary functions. Plotting it, we see that $H(y)\simeq\big{(}2-y\big{)}^{4}/\big{(}2(1-y)+e^{2a}y\big{)}^{4}$ yields a good approximation. We finally expand ${\rm sech}(x)^{e^{-4a}}\simeq 1+e^{-4a}\ln{\rm sech}(x)$, and normalize $\eta_{-1}(0)=1$, which yields the approximate ground state eigenfunction
$\eta_{-1}(x)\simeq{\rm sech}(x)^{8}\cosh(2x)^{4}\big{(}4e^{-2a}\sinh(x)^{2}+1\big{)}^{3}\big{(}1+24e^{-4a}\ln{\rm sech}(x)\big{)}/\big{(}2e^{-2a}\sinh(x)^{2}+1\big{)}^{4}$ (68)
when $a$ is large.
4.1. Time scale
The time scale of the growing perturbation mode is given by
$\tau=1/\sqrt{|\lambda_{-1}|}$ (69)
which is plotted in Fig. 9. When the sphaleron is perturbed and becomes unstable, this time scale represents the approximate lifetime of the sphaleron before the change due to instability becomes large.
Figure 9. Lifetime $\tau$
As $a\to 0$, we have $\tau\simeq\tfrac{1}{\sqrt{5}}(1+\tfrac{24}{35}a^{2})\to 0.4472$, while $\tau\simeq\tfrac{1}{4\sqrt{6}}e^{2a}\to\infty$ as $a\to\infty$.
5. Concluding remarks
The instability of sphalerons has been studied for a general quartic Klein-Gordon model with a false vacuum. Several new results are obtained.
  • The eigenfunction problem is solved exactly in terms of Heun functions.
  • In addition to an unstable ground-state mode, and the translation (zero) mode, the problem is found to have two internal vibrational modes.
  • Approximate explicit expressions in terms of elementary functions are developed for the ground state eigenvalue and eigenfunction.
An interesting question is what is the long-time behaviour of the sphaleron after it is perturbed? There are two different channels for the evolution [7]. One channel consists of the sphaleron collapsing to an oscillon, which is a long-lived oscillatory solution that has a localized profile [9, 10, 11]. This decay channel occurs when the perturbation is negative [12]. The other channel arises from a positive perturbation [12] and causes the sphaleron to increase in height until its peak reaches the value of the true vacuum, after which its width will increase, producing an expanding region of true vacuum due to conservation of energy. Its profile resembles a kink-antikink pair whose flanks steepen as they separate and accelerate.
A subsequent paper [14] will study the latter behaviour both numerically and analytically. Finally, it is also natural to study the outcome of a collision between two sphalerons before the onset of their instability [15].
Appendix A Conservation laws
Continuous symmetries of the action principle (2) give rise to conservation laws via Noether’s theorem. For a general potential, the continuous symmetries consist of time-translations, space-translations, and Lorentz boosts. These symmetries are inherited by the equation of motion (1).
Time-translations yield energy
$E[\phi]=\int_{-\infty}^{\infty}\Big{(}\tfrac{1}{2}\phi_{t}^{2}+\tfrac{1}{2}\phi_{x}^{2}+V(\phi)\Big{)}\,dx.$ (70)
Space-translations yield linear momentum
$P[\phi]=\int_{-\infty}^{\infty}\Big{(}{-}\phi_{t}\phi_{x}\Big{)}\,dx.$ (71)
Lorentz boosts yield boost momentum
$J[\phi]=\int_{-\infty}^{\infty}\Big{(}t\phi_{t}\phi_{x}+x\big{(}\tfrac{1}{2}\phi_{t}^{2}+\frac{1}{2}\phi_{x}^{2}+V(\phi)\big{)}\Big{)}\,dx.$ (72)
These three integrals are conserved for all solutions $\phi(x,t)$ with sufficient asymptotic decay as $x\to\pm\infty$.
The boost momentum can be expressed as
$J[\phi]=\chi[\phi;t]E[\phi]-tP[\phi]$ (73)
where
$\chi[\phi;t]=\frac{1}{E[\phi]}\int_{-\infty}^{\infty}x\Big{(}\tfrac{1}{2}\phi_{t}^{2}+\tfrac{1}{2}\phi_{x}^{2}+V(\phi)\Big{)}\,dx$ (74)
defines the center of energy. Conservation of $E$ and $P$ thereby implies $\frac{d}{dt}\chi=P/E$, showing that the center of energy moves at constant speed.
Acknowledgements
The author thanks Danial Saadatmand for inspiring this study. Alexandr Chernyavskiy is thanked for valuable discussions.